# Evaluating LLMs’ Inherent Multi-hop Reasoning Ability

### v4

@@ Abstract
While Large Language Models (LLMs) excel in question-answering (QA) tasks, their multi-step reasoning abilities on multiple evidence integration on Multi-hop QA tasks remain underexplored. LLMs sometimes generate answers that rely on internal memory rather than reasoning given context, which brings concerns about the evaluation quality of real reasoning abilities. The counterfactual QA task can separate internal memory from reasoning abilities, but focusing solely on final-QA performance without evaluating the multi-step reasoning process is insufficient for reporting LLMs’ real reasoning abilities. Current Multi-hop QA (MHQA) benchmarks are factual and annotated on open-source corpora such as Wikipedia, although useful for multi-step reasoning evaluation, showing limitations due to potential data contamination in LLMs pre-training stage. To address this issue, we introduce the Inherent Reasoning Evaluation (IRE) method, a novel evaluation way that jointly evaluates the LLMs’ chain-of-reasoning performance based on the first knowledge-edited counterfactual multi-hop QA data which involves editing the original Wikipedia passages, reducing data contamination risks. The IRE comprehensively assesses reasoning chains through sub-QA and final-QA evaluations. Our comparisons reveal significant performance gaps for several LLMs between Wikipedia-based benchmarks and IRE, deeming data contamination issues in existing benchmarks. We believe that the IRE benchmark will enhance and facilitate trustworthy LLM evaluations.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has demonstrated effectiveness in improving Large Language Models’ (LLMs) reasoning performance by integrating multiple pieces of evidence from a given context into the generative process, to provide accurate and relevant responses [1, 2]. However, it remains unclear whether LLMs perform reasoning, to what extent they are capable of multi-step reasoning, and how to evaluate the reasoning chain [3, 4]. Previous multi-hop QA (MHQA) benchmarks such as HotpotQA [5] and 2WikiMultihopQA [6] can be applied to evaluate LLMs’ multi-step reasoning performance on final-QA evaluation and evidence retrieval evaluation, showing limitations that final-QA and evidence retrieval are not sufficient for interpreting the multi-step reasoning ability of LLMs [7, 8, 9]. In this work, we explore whether LLMs perform the desired reasoning chain to reach the prediction and how they perform.
Previously, [10] introduced a sub-QA dataset derived from HotopotQA [5] and conducted experiments focused on sub-question reasoning. Their results reveal that strong QA models such as DFGN [11], DecompRC [12], and CogQA [13] can correctly answer multi-hop questions but often bypass the incorrect reasoning chain. This observation highlights the importance of evaluating LLMs’ multi-step reasoning abilities by assessing their performance on sub-QA. Moreover, previous MHQA benchmarks such as HotpotQA and 2WikiMultihopQA are factual and Wikipedia-based, facing the challenge of data contamination that LLMs are already pre-trained on large open-source corpora such as Wikipedia. There is a risk of data leakage in that data is memorized by the LLMs’ pre-training stage, which may inflate LLMs’ reasoning performance on MHQA tasks, complicating the assessment of their real reasoning abilities. For instance, on the top left of Figure LABEL:comparison, the question "Who was the captain of the Argentine team that was born in 1987?" is input into ChatGPT with and without the related context. ChatGPT outputs the same answer in both cases, demonstrating that it can provide the same response based on its memory regardless of the context provided.
Counterfactual QA is an effective method for disentangling LLMs’ internal memory from their reasoning abilities [14, 15, 16, 17, 18] because the given questions have answers that do not exist in the real world (counterfactual), preventing LLMs from relying solely on their internal memory to generate responses [19]. For example, the question "Which position did the coach play that serves the Chinese national basketball team and was born in 1997?" at the bottom of Figure LABEL:comparison is counterfactual because no person born in 1997 has served as a Chinese coach. ChatGPT fails to generate a correct answer from its memory without the context and can only correctly answer the question based on the given context. This observation indicates that reasoning over counterfactual contexts helps guide LLMs’ thinking based on the provided context rather than relying on pre-trained internal knowledge. Current counterfactual QA benchmarks [14, 15, 16, 17], although designed to evaluate LLMs’ real reasoning abilities, primarily assess the single-step reasoning performance of RAGs without providing a comprehensive evaluation of multi-step reasoning. To more comprehensively evaluate LLMs’ real multi-step reasoning abilities, it is vital to establish a continuous and accessible reasoning path.
To address the aforementioned limitations of factual MHQA and counterfactual QA benchmarks, we introduce IRE, which reports LLMs’ real multi-step reasoning performance on Inherent and multi-step Reasoning Evaluation that jointly computes LLMs’ reasoning performance by considering sub-QA performance as equally important as final-QA performance. To conduct experiments in a fair setting, we propose a Knowledge-edited Counterfactual MHQA dataset to comprehensively and objectively assess LLMs’ inherent multi-step reasoning ability by synthesizing counterfactual MHQA data from Wikipedia passages that LLMs may have been exposed to. Figure LABEL:comparison shows the advantages and different paradigms of our synthesized counterfactual multi-step reasoning evaluation method as well as single-step reasoning evaluation with previous Wikipedia-based MHQA evaluation, such as HotpotQA [5], 2WikiMultihopQA [6] and MuSiQue [20].
@@
We conduct massive experiments and use IRE to comprehensively and objectively report LLMs’ real reasoning ability based on equal amounts of data from previous Wikipedia-based MHQA benchmarks as the control groups and our synthesized counterfactual MHQA data. The experiment results reveal two types of inflated performance in LLMs: 1) an obvious performance gap between Wikipedia-based MHQA benchmarks and our synthesized data; 2) inflated performance due to a low proportion of correct reasoning chains, with GPT-4 achieving only 36.3% correct reasoning chains across the entire dataset. Additionally, we observe that incorporating sub-questions into the prompt as part of the reasoning chain is a more efficient approach for improving model performance. This finding underscores the significance of this method as a future research direction for enhancing LLMs’ reasoning abilities. To the best of our knowledge, we are the first to introduce counterfactual into the evaluation of the multi-step reasoning ability of LLMs, finding that there is a significant performance gap between LLMs’ performance on factual Wikipedia-based MHQA data and our synthesized counterfactual MHQA data. Furthermore, we reveal the risk of data contamination and the urgency of evaluating LLMs’ real ability in a different paradigm. The randomly selected Wikipedia-based MHQA data and our synthesized counterfactual MHQA data are available at https://anonymous.4open.science/r/LLM_inherent_multi_step_eval-3818/.
@@ 2 Related Work
Retrieval Augmented Generation RAG improves LLM’s response [21] and also mitigates the occurrence of hallucinations, thereby enhancing the models’ credibility [22]. As demonstrated by [23], designs a RAG system for multi-hop question answering and claim verification tasks. These tasks require the extraction of evidence from two or more documents to produce a correct answer. [24] proposes a Multi-hop RAG benchmark, which consists of a large collection of multi-hop queries, ground-truth answers, and the corresponding supporting evidence. Multi-hop RAG requires LLM to reason and answer multi-hop queries given the evidence. However, LLMs’ memorized knowledge sometimes conflicts with the given context, emphasizing the importance of correcting LLMs’ generations with new facts. [14, 15, 16, 17, 18] propose counterfactual QA benchmarks to separate LLMs’ parametrical knowledge (internal) and contextual knowledge (outer) that fix LLMs to reasoning on the given context strictly by editing the contextual information or prompts. Previous works motivate us to explore LLMs’ real reasoning ability by reasoning in counterfactual contexts. However, counterfactual QA datasets still only assess final QA performance and lack reasoning process evaluation.
@@
Multi-hop QA Datasets Multi-hop QA requires more than one reasoning step in multiple paragraphs to answer a question [25, 5, 6, 20]. Notably, [10] introduced a human-validated sub-question dataset derived from the HotpotQA dataset [5], undertaking a detailed investigation of the models’ capabilities to reason through sub-questions. Their findings revealed that notable models like DFGN [26], DecompRC [12], and CogQA [13] exhibit deficiencies in resolving sub-questions, even though they may successfully address the overarching multi-hop question. Moreover, Wikipedia-based MHQA datasets face the challenge of data contamination that hard to objectively and truthfully evaluate the reasoning ability of LLMs. Data contamination, i.e., the presence of test data from downstream tasks in the training data of large language models (LLMs), is a major issue in measuring LLMs’ real performance on other tasks. For example, HotpotQA [5], 2WikiMultihopQA [6], and MuSiQue [20] can be applied to evaluate the multi-step reasoning performance of LLMs. Typically, evaluating LLMs in MHQA datasets involves using RAG to retrieve and reason over context with a single step of retrieval. However, single-step retrieval can result in insufficient context retrieval for complex questions, as it provides a limited scope of information [2]. [4] proposes a framework that defines good reasoning chains in Correctness and Informativeness to illustrate whether the previous reasoning step could help the current reasoning step and the final answer. Such a framework, although useful, still faces the challenge of data contamination.
@@
Benchmarking Data Leakage [27] surveys the recent works on detecting data contamination and releases a python library named LLMSanitize that implements major contamination detection methods. A handful of recent studies have provided several strategies, methods, and benchmarks for detecting contamination without the need to access pre-training data [28, 29, 30, 31]. However, these data contamination detection benchmarks are required to dynamically update because of the development of LLMs and the expansion of pertaining data. Dynamical maintenance is time-consuming and effortless, while our proposed benchmark IRE, based on the knowledge edition, is fixed and maintains the cleanness of the test data. To statically and quantitatively detect the LLMs’ data contamination extent, [32] proposes a detection pipeline by computing perplexity and N-gram accuracy to evaluate potential data leakages. [33] designs a new Grade School Math 1000 (GSM1k) to mirror the GSM8k benchmark [34] and evaluate LLMs’ mathematical reasoning ability. Clean-Eval [35], a benchmark to assess the inflated performance of LLMs. The experimental results show a drop in performance for GPT-3.5 and Llama-2 in the Clean-Eval data, deeming the data contamination problem. [36] proposes a detection pipeline with the help of perplexity and N-gram accuracy gap between the previous MathQA dataset and synthesized data, pinpointing the potential data leakage problem. The measurement illustrates that underscores the urgency for an evaluation paradigm shift in how we approach the development and evaluation of LLMs. To properly and objectively investigate multi-step reasoning abilities, it is crucial to disentangle LLMs’ inherent memory from their reasoning capabilities. [37] introduces a novel benchmark for counterfactual QA task by editing six common reasoning schemes in the real world, while still lacking muti-step reasoning evaluation for LLMs.
3 Inherent Reasoning Evaluation
@@
In this section, we describe the pipeline of how to synthesize counterfactual MHQA data, and our novel evaluation method. As shown in Figure LABEL:Annotation_Framework, the framework has three main parts: 1) we first ask LLM such as GPT-4 to act as a passage annotator that rewrites passages from Wikipedia  [^1] https://www.wikipedia.org/ with human evaluation and feedback; 2) LLM’s automatic counterfactual QA pairs annotation with human evaluation and feedback; 3) After getting the synthesized counterfactual MHQA data, we utilize IRE to evaluate several strong LLMs to report the real multi-step reasoning performance.
3.1 Data Construction for IRE
3.1.1 Data Collection and Passage Rewriting
@@
We first randomly select 300 Wikipedia passages as the raw context. Inspired by recent studies on LLM’s ability to aid human annotation [38, 39] and the counterfactual data augmentation framework [40]. We design a pipeline for automatically annotating Wikipedia passages into counterfactual passages. Given a raw Wikipedia passage, LLMs are required to act as a passage annotator to do the named entity, noun phrase, and synonym replacement. Then we translate the replaced text into Chinese and finally back translation into English, e.g., the words in red of the annotated passage in figure LABEL:Annotation_Framework are the replaced named entities, noun phrases, and synonyms, and the new counterfactual passage is rewritten from the original Wikipedia passage. After obtaining the new and counterfactual passage, human experts manually check the data quality by 1) grammar issue, checking whether the annotated passages have grammar issues or not; and 2) making sure the annotated passage is new to LLMs. Since LLMs are pre-trained on the open-source corpus collected from the Internet, we manually search the key information of the annotated passage on the Internet, e.g. the characters, times, events, causes, processes, and results.
3.1.2 QA-pair Annotation and Checking
@@
Secondly, we also utilize LLMs to generate new multi-hop questions to fit the rewritten passages. To make sure the generated QA pairs are correct and related to the given passages, we also check the QA pairs from two perspectives: 1) Grammar issue; 2) Answerabilities of the generated question, whether the question is related to the passage and make sure the answer can be reasoned from the passages based in the generated questions. To evaluate LLMs’ performance on different complexity of multi-hop questions, for each passage, we annotate 3 complex questions: one for 2-hop, one for 3-hop, and one for 4-hop questions. LLMs are also required to generate multi-hop questions along with the corresponding sub-questions for reasoning chain evaluation. For example, at the middle part of figure LABEL:Annotation_Framework is the re-annotated context, newly generated multi-hop question, sub-questions, and intermediate answers. The annotated answers are all with a short answer span that follows the settings of HotpotQA, 2WkimultihopQA, and MuSiQue [5, 6, 20]. Here we use EM and F1 scores to measure the LLMs’ output. After getting the rewritten passages with corresponding QA pairs and sub-qa pairs, we also check the logic of the whole passage and QA pairs. We expect that the passage is coherent and the answers can be reasoned from the passage based on the given questions. The prompts of the data annotation and more annotated examples are shown in Appendix A.
3.1.3 Dataset Analysis and Statistics
@@
Table 1 shows the statistics of our dataset. We re-annotated 300 unique Wikipedia passages, with three multi-hop questions (one 2-hop, one 3-hop, and one 4-hop for each passage), a total of 3600 unique QA pairs including 900 multi-hop questions, and 2700 sub-questions with corresponding answers. Among them, the 2700 sub-questions are the decompositions of the 900 multi-hop questions. Following the settings of previous LLM evaluation benchmarks [41, 42], we treat the total of 3600 QA pairs as the test set. Followed by benchmarks such as HotpotQA [5], 2WikiMulthopQA [6], we propose a taxonomy on fine-grained question types and examples commonly used in multi-hop QA illustrated in table 8. In HotpotQA, 2WikimultihopQA, and MuSiQue datasets, the multi-hop questions are usually two types of questions, or the combination of the two types of questions, Bridge and Comparison: Bridge question is required to find the bridge entity that connects the sub-questions, while Comparison question is a type of question that compares two or more entities for the parallel sub-questions. Here we also focus on these two types and annotate multi-hop questions.
3.1.4 Human Agreement on Data Annotation
@@
After reviewing and checking the annotated data manually, we sample 300 instances randomly from our synthesized data (100 for 2-hop, 100 for 3-hop, and 100 for 4-hop) and assign each instance to two reviewers (as expert human annotators). We ask them to select and discard the annotated data that badly follow the given guidelines. The experts first checked the grammar of the annotated questions, sub-questions, sub-answers, and answers. Then they search the knowledge of the annotated passages on the Internet to see whether thaose corpus has appeared or not. According to our statistics, the agreement rate between the annotators in the randomly selected IRE samples is 94% and the human agreement rates are 96%, 92%, and 95% in the 2-hop, 3-hop, and 4-hop datasets, respectively. This suggests that our synthesized instances reflect good data quality on annotation guideline following and achieve high human agreement among expert annotators. To quantitatively illustrate data quality, we also utilize GPT-4 to assign scores to each selected data from two perspectives: correctness and informativeness. Each data is assigned with 1 or 0, which means correct or incorrect, informative or not informative. Correctness indicates whether the answer can be reasoned from the given question and context, while informativeness means whether the QA pairs and context are related or not.
@@
2 hop dataset 3 hop dataset 4 hop dataset Whole data property Value 2 hop QA pairs 300 3 hop QA pairs 300 4 hop QA pairs 300 Unique Passages 300 Sub-QA pairs 600 Sub-QA pairs 900 Sub-QA pairs 1200 Total QA pairs 3600 Correctness 94 Correctness 93 Correctness 92 Sentences per data (Avg) 38.42 Informativeness 85 Informativeness 86 Informativeness 82 Inter-annotator Agreement 94%
Table 1: The statistics of each subset of our synthesized data and the whole dataset.
Benchmarks Data Type Data Source Task Reasoning Chain HotpotQA Factual Wikipedia Multi-hop/final-QA 2WikiMultihopQA Factual Wikipedia Multi-hop/final-QA MusiQue Factual Wikipedia Multi-hop/final-QA DisentQA Counterfactual & Factual Natural Questions Single-hop/final-QA IfQA Counterfactual Crowdsourcing Open Domain/final-QA Ours Counterfactual Rewriting Wikipedia Multi-hop/final-QA/sub-qa
Table 2: Differences between our synthesized data with previous factual and counterfactual QA evaluation benchmarks.
3.2 Evaluation Method
The multi-hop QA evaluation is referred to as finding answers for complex questions that require reasoning multiple times from given passages. We employ three representative QA evaluation methods to assess the correctness of LLM-generated MHQA responses: sub-question answering evaluation, reasoning chain evaluation, and the joint performance of sub-qa and MHQA.

### v5

# Cofca: A Step-Wise Counterfactual Multi-hop QA benchmark

@@ Abstract
While Large Language Models (LLMs) excel in question-answering (QA) tasks, their real reasoning abilities on multiple evidence retrieval and integration on Multi-hop QA tasks remain less explored. Firstly, LLMs sometimes generate answers that rely on internal memory rather than retrieving evidence and reasoning in the given context, which brings concerns about the evaluation quality of real reasoning abilities. Although previous counterfactual QA benchmarks can separate the internal memory of LLMs, they focus solely on final QA performance, which is insufficient for reporting LLMs’ real reasoning abilities. Because LLMs are expected to engage in intricate reasoning processes that involve evidence retrieval and answering a series of sub-questions from given passages. Moreover, current factual Multi-hop QA (MHQA) benchmarks are annotated on open-source corpora such as Wikipedia, although useful for multi-step reasoning evaluation, they show limitations due to the potential data contamination in LLMs’ pre-training stage. To address these issues, we introduce a Step-wise Counterfactual benchmark (CofCA), a novel evaluation benchmark consisting of factual data and counterfactual data that reveals LLMs’ real reasoning abilities on multi-step reasoning and reasoning chain evaluation. Our experimental results reveal a significant performance gap of several LLMs between Wikipedia-based factual data and counterfactual data, deeming data contamination issues in existing benchmarks. Moreover, we observe that LLMs usually bypass the correct reasoning chain, showing an inflated multi-step reasoning performance. We believe that our CofCA benchmark will enhance and facilitate the evaluations of trustworthy LLMs.
@@ 1 Introduction
Retrieval-augmented generation (RAG) (Chen et al., 2023; Ma et al., 2023; Gao et al., 2023b), has garnered increasing research attention (Tang & Yang, 2024; Poliakov & Shvai, 2024). As shown in Figure 1a, given a query, a RAG system first retrieves pieces of evidence from a set of relevant passages, and then generates a response by taking the passages as additional context to LLMs. While most existing work concentrates on the setting where the answer can be derived from a single retrieved passage, there are situations where LLMs are required to integrate information from multiple resources as evidence to infer the answer. As shown in Figure 1b, given the query, the single retrieved passages cannot offer enough information to generate correct answers. Instead, the evidence “Earth is the third planet” and “Mars is after Earth” are required to be integrated for deciding “The fourth planet of the sun”.
Recently, a line of work (Wang et al., 2023b; Staniszewski et al., 2023; Trivedi et al., 2022) has investigated the evidence integration capabilities of LLMs combined with RAG, using existing Multi-hop QA (MHQA) datasets such as HotpotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). However, such work focuses only on the correctness of the final answer, without evaluating the reasoning process itself. It is thus difficult to understand whether shortcut reasoning exists (Tang et al., 2021; Ho et al., 2022; Yang et al., 2022) and whether such reasoning is robust across out-of-distribution (OOD) tasks. Existing research (Tang et al., 2021) shows that strong QA models such as DFGN (Xiao et al., 2019), DecompRC (Min et al., 2019), and CogQA (Ding et al., 2019) can correctly answer multi-hop questions but bypass the correct reasoning chain. However, no existing research has evaluated the performance of LLMs to the same problem.
Figure 1: Differences between (a) and(b). The words in red are the pieces of evidence for the given questions. The differences between factual MHQA (c) and counterfactual MHQA (d). c: we input a factual MHQA with and without context into ChatGPT. ChatGPT could output the correct answer based on its internal memory regardless of the context. d: When inputting a counterfactual QA, where the passage is rewritten from the existing factual passage in c (words in the red), ChatGPT cannot rely on its memory and must reason on the given context, deeming that counterfactual QA can decouple LLMs’ internal memory and reasoning abilities.
We aim to address the above issues by making a systematic evaluation of the LLMs’ reasoning process. To this end, an existing sub-question style benchmark (Tang et al., 2021) derived from HotpotQA Yang et al. (2018) can be useful, in which LLMs are required to answer a series of sub-questions to reach the target rather than giving a final answer solely. However, existing datasets cannot be directly used for reporting LLM real performance, due to the risk of data leakage in the pretraining stage (L’opez et al., 2024; Zhou et al., 2023a; Fu et al., 2023). As shown in Figure 1c, for instance, given the question “Who was the captain of Argentine team that was born in 1987?”, ChatGPT might give the correct answer even without a context, since relevant knowledge is learned by the model during the pre-training stage. Therefore, a sub-question benchmark that avoids the negative influence of data leakage is needed. To this purpose, we consider the counterfactual evaluation method (Li et al., 2023; Neeman et al., 2023; Zhou et al., 2023b; Wu et al., 2024; Yu et al., 2023). We make modifications to existing knowledge in Wikipedia, deriving new passages and ensuring that knowledge in these passages is non-existent in the real world, thus those answers have little chance of being inherent in LLMs. For example, in Figure 1d, given the question “Which position did the coach play that serves Chinese national basketball team and born in 1997?”, LLMs cannot rely on their memories and must answer the first sub-question “Which coach serves as Chinese national basketball team and born in 1997?”, before finding the final answer by asking a second sub-question “Which position did Alice Brown play?”. The passage in Figure 1d, is modified from the passage in Figure 1c. Specifically, we generate counterfactual data to Wikipedia knowledge by replacing key information (named entities, noun phrases, and synonyms) and then paraphrasing (back translation).
@@
We call our dataset CofCA, a step-wise and Counterfactual MHQA benchmark, which consists of counterfactual passages with corresponding multi-hop QA pairs and sub-QA pairs, and an equal amount of factual MHQA data from HotpotQA, 2WikiMultihopQA and MuSiQue as the control group. Using CofCA, we evaluate and report LLMs’ real reasoning ability. Experimental results show two types of inflated performance in LLMs: 1) an obvious performance gap between Wikipedia-based factual MHQA benchmarks and counterfactual MHQA data; 2) inflated performance due to a low proportion of correct reasoning chains as well as a high proportion of incorrect reasoning chains, with GPT-4 achieving only 36.3% correct reasoning chains across the entire dataset. Additionally, we observe that incorporating sub-questions into the prompt as part of the reasoning chain is a more efficient approach for improving LLMs’ performance.
To the best of our knowledge, we are the first to introduce counterfactual data into the evaluation of the multi-step reasoning ability of LLMs, finding that there is a significant performance gap between LLMs’ performance on factual data and counterfactual data. Furthermore, the reasoning abilities of LLMs are exaggerated because of the poor reasoning chain performance. The Whole CofCA data are available at https://anonymous.4open.science/r/LLM_inherent_multi_step_eval-3818/.
@@ 2 Related Work
Retrieval Augmented Generation (RAG) improves LLM’s response (Borgeaud et al., 2021) and also mitigates the occurrence of hallucinations, thereby enhancing the models’ credibility (Gao et al., 2023a). As demonstrated by Khattab et al. (2021), designs a RAG system for MHQA and claim verification tasks. These tasks require the extraction of evidence from two or more documents to produce a correct answer. Tang & Yang (2024) propose a Multi-hop RAG benchmark, which consists of a large collection of multi-hop queries, ground-truth answers, and the corresponding supporting evidence. Multi-hop RAG requires LLM to reason and answer multi-hop queries given the evidence. However, LLMs’ memorized knowledge sometimes conflicts with the given context, emphasizing the importance of correcting LLMs’ generations with new facts. Li et al. (2023); Neeman et al. (2023); Zhou et al. (2023b); Wu et al. (2024); Yu et al. (2023) propose counterfactual QA benchmarks to separate LLMs’ parametrical knowledge (internal) and contextual knowledge (outer) that fix LLMs to reasoning on the given context strictly by editing the contextual information or prompts. Previous work motivates us to explore LLMs’ real reasoning ability by reasoning in counterfactual contexts. However, counterfactual QA datasets still only assess final QA performance and lack reasoning process evaluation.
@@
Multi-hop QA Datasets Multi-hop QA requires more than one reasoning step in multiple paragraphs to answer a question (Dua et al., 2019; Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2021). Notably, Tang et al. (2021) introduce a human-validated sub-question dataset derived from the HotpotQA dataset (Yang et al., 2018), undertaking a detailed investigation of the models’ capabilities to reason through sub-questions. Their findings revealed that notable models like DFGN (Xiao et al., 2019), DecompRC (Min et al., 2019), and CogQA (Ding et al., 2019) exhibit deficiencies in resolving sub-questions, even though they may successfully address the overarching multi-hop question. Moreover, Wikipedia-based MHQA datasets face the challenge of data contamination that hard to objectively and truthfully evaluate the reasoning ability of LLMs. Data contamination, i.e., the presence of test data from downstream tasks in the training data of large language models (LLMs), is a major issue in measuring LLMs’ real performance on other tasks. For example, HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021) can be applied to evaluate the multi-step reasoning performance of LLMs. Typically, evaluating LLMs in MHQA datasets involves using RAG to retrieve and reason over context with a single step of retrieval. However, single-step retrieval can result in insufficient context retrieval for complex questions, as it provides a limited scope of information (Gao et al., 2023b). Prasad et al. (2023) proposes a framework that defines good reasoning chains in Correctness and Informativeness to illustrate whether the previous reasoning step could help the current reasoning step and the final answer.
@@
Benchmarking Data Leakage A handful of recent studies have provided several strategies, methods, and benchmarks for detecting contamination without the need to access pre-training data (Shi et al., 2023; Roberts et al., 2023; Golchin & Surdeanu, 2023; Zhu et al., 2023a). Ravaut et al. (2024) surveys recent work on detecting data contamination and releases a python library named LLMSanitize that implements major contamination detection methods. However, these data contamination detection benchmarks are required to dynamically update because of the development of LLMs and the expansion of pertaining data. Dynamical maintenance is time-consuming and effortless, while our proposed benchmark CofCA, based on the knowledge edition, is fixed and maintains the cleanness of the test data. To statistically and quantitatively detect the LLMs’ data contamination extent, Xu et al. (2024b) proposes a detection pipeline by computing perplexity and N-gram accuracy to evaluate potential data leakages. Zhang et al. (2024) designs a new Grade School Math 1000 (GSM1k) to mirror the GSM8k benchmark (Cobbe et al., 2021) and evaluate LLMs’ mathematical reasoning ability. Clean-Eval (Zhu et al., 2023b), a benchmark to assess the inflated performance of LLMs. The experimental results show a drop in performance for GPT-3.5 and Llama-2 in the Clean-Eval data, deeming the data contamination problem. Xu et al. (2024a) proposes a detection pipeline with the help of perplexity and N-gram accuracy gap between the previous MathQA dataset and synthesized data, pinpointing the potential data leakage problem. The measurement illustrates that underscores the urgency for an evaluation paradigm shift in how we approach the development and evaluation of LLMs. To properly and objectively investigate multi-step reasoning abilities, it is crucial to disentangle LLMs’ inherent memory from their reasoning capabilities. Hua et al. (2024) introduces a novel benchmark for counterfactual QA task by editing six common reasoning schemes in the real world, while still lacking muti-step reasoning evaluation for LLMs. Our work is different from such works in two main aspects: 1) a comprehensive reasoning chain evaluation; and 2) Counterfactual Question Answering that reduces the risk of data contamination.
3 Multi-hop Reasoning Evaluation
@@
In this section, we design a human-in-the-loop annotation framework and our novel evaluation method. As shown in Figure LABEL:Annotation_Framework, the framework consists of three components: 1) we first ask LLM such as GPT-4 to act as a passage annotator that rewrites passages from Wikipedia with human evaluation and feedback; 2) LLM’s automatic counterfactual QA pairs annotation with human evaluation and feedback; 3) After obtaining the synthesized counterfactual MHQA data, we use counterfactual multi-hop reasoning to evaluate several strong LLMs to report the real multi-step reasoning performance.
3.1 Data Construction
Data Collection and Passage Rewriting
@@ 3.1 Data Construction
We randomly select 300 Wikipedia passages from Huggingface [^1] https://huggingface.co/datasets/lucadiliello/english_wikipedia as the raw context. Inspired by recent studies on LLM’s ability to aid human annotation (Bartolo et al., 2021; Törnberg, 2023), we design a pipeline for automatically annotating Wikipedia passages into counterfactual passages. Given a raw Wikipedia passage, LLMs are required to act as a passage annotator to do the named entity, noun phrase, and synonym replacement. Since LLMs are pre-trained on the open-source corpus collected from the Internet, we manually search the key information of the annotated passage on the Internet, e.g. the characters, times, events, causes, processes, and results.
After the replacement stage, we translate the replaced text into Chinese and finally back translation into English, e.g., the words in red of the annotated passage in figure LABEL:Annotation_Framework are the replaced named entities, noun phrases, and synonyms, and the new counterfactual passage is rewritten from the original Wikipedia passage. Upon acquiring the new and counterfactual passages, human experts conduct a manual evaluation of data quality by focusing on two key aspects: 1) assessing the grammatical integrity of the annotated passages to ensure there are no grammatical errors; and 2) verifying that the annotated passages are new and counterfactual to Large Language Models (LLMs).
QA-pair Annotation and Checking
@@
We use LLMs to generate new multi-hop questions to fit the rewritten passages. To make sure the generated QA pairs are correct and related to the given passages, we check the QA pairs from two perspectives: 1) Grammar issue; 2) Answerabilities of the generated question, whether the question is related to the passage and make sure the answer can be reasoned from the passages based in the generated questions. To evaluate LLMs’ performance on different complexity of multi-hop questions, for each passage, we annotate 3 complex questions: one for 2-hop, one for 3-hop, and one for 4-hop questions. LLMs are required to generate multi-hop questions along with the corresponding sub-questions for reasoning chain evaluation. For example, at the middle part of Figure LABEL:Annotation_Framework illustrates the re-annotated context, newly generated multi-hop question, sub-questions, and intermediate answers. The annotated answers are all with a short answer span that follows the settings of HotpotQA, 2WkimultihopQA, and MuSiQue (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2021). Here we use EM and F1 scores to measure the LLMs’ output. After obtaining the rewritten passages with corresponding QA pairs and sub-qa pairs, we also check the logic of the whole passage and QA pairs. We expect that the passage is coherent and the answers can be reasoned from the passage based on the given questions. The prompts of the data annotation and more annotated examples are shown in Appendix A.
Dataset Analysis and Statistics
@@
Table 1 shows the statistics of our dataset. We re-annotated 300 unique Wikipedia passages, with three multi-hop questions (one 2-hop, one 3-hop, and one 4-hop for each passage), a total of 3600 unique QA pairs including 900 multi-hop questions, and 2700 sub-questions with corresponding answers. Besides, we also randomly select 900 factual MHQA data as the control group (300 from HotpotQA, 300 from 2WikiMultihopQA, and 300 from MuSiQue). Following the settings of previous LLM evaluation benchmarks (Wang et al., 2023a; 2024), we treat the total of 1800 data as the test set. Following benchmarks such as HotpotQA (Yang et al., 2018), 2WikiMulthopQA (Ho et al., 2020), we propose a taxonomy on fine-grained question types and examples commonly used in multi-hop QA illustrated in Table 11. In HotpotQA, 2WikimultihopQA, and MuSiQue datasets, the multi-hop questions consists of two types of questions, or the combination of the two types of questions, Bridge and Comparison: Bridge question is required to find the bridge entity that connects the sub-questions, while Comparison question is a type of question that compares two or more entities for the parallel sub-questions. We focus on these two types of questions for annotating multi-hop questions.
Inter Human Agreement
Following a thorough manual inspection of the annotated data, we randomly select 300 samples, dividing them equally among 2-hop, 3-hop, and 4-hop instances, with each sample being evaluated by two expert reviewers. These reviewers are tasked with identifying and eliminating any data that significantly deviates from the established guidelines. Initially, the experts focus on assessing the grammatical correctness of the questions, sub-questions, sub-answers, and answers within the annotations. Subsequently, they verify the originality of the content in the annotated passages by checking for their presence on the Internet.
@@
According to our statistics, the agreement rate between the annotators in the randomly selected CofCA samples is 94% and the human agreement rates are 96%, 92%, and 95% in the 2-hop, 3-hop, and 4-hop datasets, respectively. This suggests that our synthesized instances reflect good data quality on annotation guideline following and achieve high human agreement among expert annotators. To quantitatively illustrate data quality, we also utilize GPT-4 to assign scores to each selected data from two perspectives: correctness and informativeness. Each data is assigned with 1 or 0, which means correct or incorrect, informative or not informative. Correctness indicates whether the answer can be reasoned from the given question and context, while informativeness means whether the QA pairs and context are related or not.
@@
2 hop dataset 3 hop dataset 4 hop dataset Whole data property Value 2 hop QA pairs 300 3 hop QA pairs 300 4 hop QA pairs 300 Unique Passages 1800 Sub-QA pairs 600 Sub-QA pairs 900 Sub-QA pairs 1200 Total QA pairs 4500 Correctness 94 Correctness 93 Correctness 92 Sentences per data (Avg) 38.42 Informativeness 85 Informativeness 86 Informativeness 82 Inter-annotator Agreement 94%
Table 1: Statistics of our CofCA dataset.
Benchmarks Data Type Data Source Task Reasoning Chain HotpotQA Factual Wikipedia Multi-hop/final-QA 2WikiMultihopQA Factual Wikipedia Multi-hop/final-QA MusiQue Factual Wikipedia Multi-hop/final-QA DisentQA Counterfactual & Factual Natural Questions Single-hop/final-QA IfQA Counterfactual & Factual Crowdsourcing Open Domain/final-QA CofCA(Ours) Counterfactual & Factual Rewriting Wikipedia Multi-hop/final-QA/sub-qa
Table 2: Differences between our CofCA with previous factual and counterfactual QA benchmarks.
3.2 Evaluation Metrics
Given a set of input documents, we employ three representative QA evaluation methods to assess the correctness of LLM-generated MHQA responses: sub-question answering evaluation, reasoning chain evaluation, and the joint performance of sub-qa and MHQA.